# Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos

### v2

@@ Abstract
We use FIRE-2 cosmological zoom-in hydrodynamic simulations to investigate the co-evolution between Milky Way-size galaxies and their host dark matter halos. We find that the formation of these galaxies follows a two-phase pattern, with an early phase featured by hot dynamics, bulge-dominated structure and bursty star formation, and a later phase featured by cold dynamics, disk-dominated structure and steady star formation. The transition times of these galaxy properties are tightly correlated with the time when the host halo transits from fast to slow accretion, suggesting that the two-phase assembly of halos drive the two-phase formation of galaxies. The physical origin of dynamical hotness can be summarized into two modes of star formation: a scattered mode in which stars form at large radii within cold gas streams associated with fast assembly of halos, and a concentrated mode in which stars form at small radii through violent fragmentation from globally self-gravitated gas when halo assembly is about to slow down. Cold gaseous and stellar disks can form only when the conditions of the two modes are removed by the stall of halo assembly and the reduction of gas by feedback processes. The two modes of star formation leave distinct imprints on the structural properties of high-redshift galaxies, providing implications to be tested by JWST and future observations.
@@
pubyear: 2025pagerange: Two-phase formation of galaxies: the coevolution between galaxies and dark matter halosA
@@ 1 Introduction
Galaxies exhibit diverse morphologies ranging from those dominated by bulges to those dominated by disks – known as the Hubble sequence (e.g. Hubble, 1926, 1936; Sandage, 1961; Kormendy, 1979; Conselice, 2014). Observations with resolved kinematic measurements suggest that the morphologies of galaxies are closely related to the motion of stars, with bulges being populated mainly by stars with random orbits and disks with ordered rotation (e.g. Cappellari, 2016; Zhu et al., 2018). Historically, these two types of motion were inferred to reflect the evolution stages of galaxies, with ordered rotation being the early products of galaxy formation owing to the collapse of gas with initial angular momentum in halos, and random motion being generated by later ‘heating’ processes such as mergers (e.g. Hernquist, 1992; Shen et al., 2003; Lacey et al., 2016). Such a picture of galaxy formation and morphological transformation has been widely adopted in theoretical models and achieved great success in reproducing the observed properties of the galaxy population observed in the local Universe, such as the size-mass relations (Shen et al., 2003; van Dokkum et al., 2008, 2015; Bezanson et al., 2009; Newman et al., 2012; Graham, 2013; van der Wel et al., 2014a; Somerville et al., 2018; Miller et al., 2019) and the bimodality in the color-stellar mass relation (Strateva et al., 2001; Baldry et al., 2004).
@@
Along this line, Mo et al. (2024, hereafter MCW24) proposed a ‘two-phase’ scenario of galaxy formation, in which the two morphological components of galaxies, bulge and disk, naturally arise from cosmological structure formation. This scenario is motivated by the fact that the assembly of individual CDM halos follows a general pattern consisting of two phases: an early phase of fast accretion and a later phase of slow accretion – a discovery made decades ago in N-body simulations (Zhao et al. 2003; see also Wechsler et al. 2002; Zhao et al. 2009; Correa et al. 2015). During the fast phase of a halo, frequent perturbation of gas is produced by the rapid change of gravitational potential associated with the fast accretion, and is subsequently amplified by shocks in supersonic turbulence generated by rapid gas inflow. Dense sub-clouds are produced by the fragmentation of the collapsing gas. The rich amount of gas and the strong turbulence brought in by the fast accretion imply that the fragmentation starts once the gas becomes self-gravitating and that the formation of dense sub-clouds proceeds rapidly before the gas can dissipate to form a dynamically cold disk. A galaxy formed in this phase is thus manifested as a self-gravitating gas cloud (SGC) filled with sub-clouds with random orbits, with star clusters born out of the sub-clouds inheriting their hot dynamics. Dynamically cold gaseous and stellar disk can form only after the host halo transits to the slow phase and the gas content is reduced by feedback processes. Implemented semi-analytically into subhalo mergers trees, this scenario can reproduce a wide range of observables, and coherently predict the co-evolution among halos, galaxies, supermassive black holes (SMBHs), and globular clusters (GCs) from the seeding era to the present day (see Mo et al., 2024; Chen et al., 2024, 2025b, 2025a, for details).
@@
One important prediction of the two-phase scenario is a counterintuitive order in the build-up of the Hubble sequence: the dynamically hot (bulge) component of a galaxy is actually built first, and followed by the dynamically cold (disk) component, with a transition corresponding to that in halo assembly from fast to slow. This scenario seems to be in conflict with the historical picture of galaxy formation and morphological transformation mentioned above, but becomes increasingly supported by recent hydrodynamic simulations (Dekel et al., 2009; Pillepich et al., 2019; Hopkins et al., 2018; Jung et al., 2025) and observations (Kassin et al., 2012; Swinbank et al., 2012; Wisnioski et al., 2015; Simons et al., 2017; Price et al., 2020). However, many questions remain unanswered. For example, it is unclear whether the fast accretion can indeed sustain a gas-rich and turbulent SGC, whether sub-clouds can actually form in the SGC quickly before settling into a rotation-supported disk, and whether feedback processes are strong enough to reduce the gas content to allow the formation of a stable disk. Given the challenges in obtaining a complete sample of high-redshift galaxies to perform dynamical decomposition, it is still difficult to answer these questions directly using observational data. On the other hand, modern zoom-in hydro simulations with cosmological initial conditions are able to resolve key processes involved in the formation of individual galaxies, thus providing opportunities to test the viability of the two-phase model.
@@
In this paper, we use a set of such simulations performed as a part of the Feedback In Realistic Environments (FIRE) project (Hopkins et al., 2018) to analyze the co-evolution between galaxies and dark matter halos, focusing on the aforementioned questions. This paper is organized as follows. In §2, we describe the simulation and the method to quantify halos and galaxies. In §3, we verify the two-phase growth pattern of galaxies and its relation to the two-phase assembly of halos. In §4, we analyze in detail the physical processes that drive the two-phase growth of galaxies. In §5, we provide implications for observations in a search for evidence induced by these processes. Finally, we summarize and discuss our findings in §6.
@@ 2 The Simulation
Table 1: The sample of FIRE-2 galaxies used in this paper. The properties listed in the columns are: (i) the name of the galaxy, with the reference first introducing the galaxy indicated in the parentheses (A, Garrison-Kimmel et al. 2019a; B, Garrison-Kimmel et al. 2019b; C, Garrison-Kimmel et al. 2017; D, Wetzel et al. 2016; E, Hopkins et al. 2018; F, Samuel et al. 2020); (ii) $M_{\mathrm{vir},\mathrm{dm}}$, the mass of dark matter enclosed in the virial radius of the host halo at $z=0$; (iii) $M_{*}$, stellar mass of the galaxy; (iv) $R_{50,*}$, half-stellar-mass radius using all star particles linked to the galaxy; (v) $R_{\rm 50,bulge}$, half-stellar-mass radius using only bulge stars; (vi) $R_{90,*}$, the radius enclosing 90% of the stellar mass, using all star particles linked to the galaxy; (vii) $t_{\mathcal{B}}$, the transition time of burstiness; (viii) $t^{(\mathrm{start})}_{\mathcal{H}}$$t^{(\mathrm{end})}_{\mathcal{H}}$, the transitional interval of hotness; (ix) $^{(\mathrm{start})}_{\mathrm{thin}}tt$$^{(\mathrm{end})}_{\mathrm{thin}}$, the transitional interval of thin-disk fraction; (x) $t_{\rm bulge,1/2}$, the half-mass formation time of stellar bulge; (xi) $t^{(\mathrm{start})}_{\gamma}$$t^{(\mathrm{end})}_{\gamma}$, the transitional interval of halo accretion rate; (xii) $t_{\rm h,1/2}$, the half-mass formation time of the host halo; (xiii) $t_{\text{h, Vpeak}}$, the epoch when the maximum circular velocity of the halo reaches the peak during the assembly history. See §2 for the details of the sample and the definitions of the properties.
@@
In this paper, we use the cosmological hydrodynamic zoom-in simulations from FIRE-2, a part of the FIRE project, performed using a multimethod gravity plus (magneto)hydrodynamics code GIZMO (Hopkins, 2015), with updated numerical methods and physical models as detailed in Hopkins et al. (2018). FIRE-2 adopted the mesh-free Lagrangian Godunov (MFM) method for hydrodynamics, which provides adaptive spatial resolution while maintaining exact conservation of mass, energy, and momentum, excellent angular momentum conservation, and accurate shock capturing. Radiative heating and cooling for gas across a temperature range of $\sim 10$$10^{10}\,{\rm K}$ are included, involving free-free, photoionization/recombination, Compton, photo-electric, metal-line, molecular, fine-structure, dust collisional, and cosmic-ray processes. Eleven elements (H, He, C, N, O, Ne, Mg, Si, S, Ca, Fe) are self-consistently tracked. Unresolved turbulent diffusion of metals in the gas is explicitly modeled (Hopkins, 2016; Su et al., 2017; Escala et al., 2018). The simulations also include photoionization and heating from a redshift-dependent, spatially uniform ultraviolet (UV) background (Faucher-Giguère et al., 2009) that reionizes the universe at $z\sim 10$. Collisionless star particles, representing ensembles of stars, are bred from molecular gas satisfying self-gravitating, self-shielding, Jeans unstable conditions and exceeding a minimum density threshold of $1000\,{\rm cm}^{-3}$ (Krumholz & Gnedin, 2011). FIRE-2 implemented distinct channels of stellar feedback, including that from OB stars, AGB mass-loss, type Ia and type II supernovae, photoelectric heating, and radiation pressure (Hopkins et al., 2018, 2020). Such feedback regulates the star formation efficiency of star-forming gas to a level of $\sim 1\%$$10\%$ per free-fall time (Faucher-Giguère et al., 2013; Hopkins, 2017; Orr et al., 2018). We note that feedback from accreting SMBHs is not included in the simulations.
@@
Figure 2: Properties of the galaxy m12f and its host halo as functions of the cosmic time ($t$) or redshift ($z$). Each panel shows a different property. (a), stellar mass in total (gray) and separately for the three kinematic components (colored), and $M_{\rm vir}$ of the host halo (black). Three vertical lines indicate the epochs at which the bulge mass reaches 20%, 50% and 80%, respectively, of its final value at $z=0$, with that corresponding to 50% denoted as $t_{\rm bulge,1/2}$. (b), the growth rate of the three kinematic components (colored). For reference, we show the accretion rate of baryons, defined as the growth rate of $M_{\rm vir}$ times the cosmic baryon fraction (black). (c), mass fractions of the three kinematic components, using young stars only. Grey vertical band indicates the transitional interval, $_{\rm tnin}^{\rm(start)}tt$$_{\rm thin}^{\rm(end)}$, of the thin-disk fraction. (d), dynamical hotness of the galaxy ($\mathcal{H}$), calculated using young stars only. Grey vertical band indicates the transitional interval, $t_{\mathcal{H}}^{\rm(start)}$$t_{\mathcal{H}}^{\rm(end)}$ of $\mathcal{H}$. (e), burstiness of star formation ($\mathcal{B}$). Vertical line indicates $t_{\mathcal{B}}$, the epoch when the star formation transits from bursty to steady (defined by $\mathcal{B}=0.2$, indicated by the dashed horizontal line). (f), cold-gas spin $\lambda_{\text{cold-gas}}$, and fractions of cold ($f_{\text{cold-gas}}$), warm and hot gas. (g), (h), virial velocity ($V_{\rm vir}$) and specific growth rate of the host halo. Grey vertical bands indicate the transitional interval, $t_{\gamma}^{\rm(start)}$$t_{\gamma}^{\rm(end)}$, of halo assembly, defined by $\gamma=3/8$0 (horizontal lines in h). In (c), (d), (g) and (h), each solid curve is obtained from the simulation, while the associated dashed curve is obtained by a parametric fitting (see §2.1 and §2.2 for details). This figure clearly demonstrate the two-phase evolution of the galaxy and its host halo. See §3 for details.
@@ 2.1 Properties of the simulated galaxies
$\epsilon$ Orbital circular of a particle (§2.1.1)
@@ 2.1.1 Stellar kinematics and decomposition
We follow the decomposition method of Abadi et al. (2003) to separate member star particles of each galaxy at a given snapshot into three components: bulge, thick disk and thin disk. The method is based on the circularity of each star particle,
@@
where $j_{z}$ is the angular momentum of that particle projected to the direction of $\bm{J}_{\rm tot}$, the net angular momentum of all stellar particles within $R_{\rm 90,*}$ at that snapshot (Gurvich et al., 2023), $j_{\rm c}(E)$ is the angular momentum of a circular orbit with the same energy ($E$) as that of the star particle. Particles with $1\leqslant\epsilon<0.2$, $0.2\leqslant\epsilon<0.8$ and $0.8\leqslant\epsilon\leqslant 1$ are categorized as bulge, thick disk and thin disk, respectively. Previous studies adopted more rigorous kinematic classification methods (e.g. Du et al., 2019, 2020; Zana et al., 2022; Liang et al., 2025), such as classifications in the plane of angular momentum versus orbital energy. However, our work focuses primarily on morphological transitions, particularly in the evolution of the significance of the disk component, which can be characterized robustly using the angular momentum alone. We therefore employ a simplified kinematic decomposition based solely on the angular momentum to define the three components. We define the direction of $\bm{J}_{\rm tot}$ as the normal direction (the $z$ axis) of the disk ($x$-$y$) plane. The projections of a simulated galaxy perpendicular to and along the $z$-axis are defined as the edge-on and face-on views of the galaxy, respectively.
@@
Fig. 1 shows an example galaxy, m12f, at $z=0$, with the three components shown separately and in total, viewed edge-on and face-on. Here, we include all particles within $R_{\rm 90,*}$. Panel (e) shows the distribution of $\epsilon$ for all stellar particles within $R_{\rm 90,*}$, weighted by their masses. Panel (f) shows the face-on surface density profile of the three components and of the total, using all particles within $R_{\rm 90,*}$. The thin disk is extended, dominates the surface density in the outer region, and follows an exponential profile at $r\gtrsim 3\,{\rm{kpc}}\approx R_{\rm 50,*}$. In the inner region ($r\lesssim 1\,{\rm{kpc}}$), however, the thin disk is truncated. This truncation is similar to that found in other simulations (Du et al. 2020, see their fig. 1; Buck et al. 2019, see their fig. 2) and in observations with spatially-resolved spectroscopy (Zhu et al. 2018, see their fig. 8 and §4), and suggests a potential caveat of using an untruncated profile to fit the entire cold-disk component. On the other hand, the thick disc is more vertically extended, reflecting its hotter dynamics with a higher degree of random-motion support for the vertical structure. Both the images (c1, c2) and the surface density profile (yellow in f) suggest the existence of an inner, concentrated component at around $1\,{\rm{kpc}}$, in alignment with the concept of slow-rotating ‘pseudo-bulge’ in the literature (Kormendy & Kennicutt, 2004). The bulge dominates the inner $0.5\,{\rm{kpc}}$, but it also has a faint extension to the outer region ($\gtrsim 10\,{\rm{kpc}}$), consistent with being a combination of the ‘classical bulge’ and the stellar halo found in the literature (Genel et al., 2015; Du et al., 2020; Zana et al., 2022). The three components have similar mass within $R_{\rm 90,*}$, as can be seen from the mass-weighted distribution of $\epsilon$ (panel e), consistent with the results of an earlier analysis of the FIRE-2 by Yu et al. (2021). All these show that FIRE-2 is able to produce realistic kinematics of MW-size galaxies, with components similar to those found in observations.
@@
To quantify the kinematics of star-forming gas within a target galaxy, we utilize in-situ young stars, defined as bound stellar particles with ages $\leq 100\,{\rm Myr}$ and birthplaces within $\leq R_{90,*}$ from the center of the galaxy. This selection is made to avoid complexity introduced by potential post-formation heating of old stars and the contamination of ex-situ stars. Hereafter, we use these in-situ young stars to define transition times in the morphological type of galaxy. As shown in Fig. 1(e), the distribution of $\epsilon$ of in-situ young stars at $z=0$ is sharply peaked around $\epsilon=1$. Most ($\approx 93\%$) of the in-situ young stars at this epoch belong to the thin disk, and only a small fraction ($\approx 1.6\%$) is categorized into the bulge component. Fig. 1(f) reinforces this conclusion by comparing the face-on surface densities of in-situ young stars (dashed lines) with those of all stars (solid lines), which shows that in-situ young stars are mostly contained in the extended thin disk.
@@
The clear transition of star formation from the dynamically hot to cold components shown in Fig. 2(c) motivates us to define a transition time to separate the early, hot phase from the late, cold phase, for each individual galaxy. To do this, we fit the fraction of young stars in the thin-disk component ($f_{\rm thin}$) as a function of the cosmic time ($t$), as shown by the blue solid curve in Fig. 2(c) for example, by the following piecewise function:

### v3

@@ Abstract
We use FIRE-2 cosmological zoom-in hydrodynamic simulations to investigate the co-evolution between Milky Way-size galaxies and their host dark matter halos. We find that the formation of these galaxies follows a two-phase pattern, with an early phase featured by hot dynamics, bulge-dominated structure and bursty star formation, and a later phase featured by cold dynamics, disk-dominated structure and steady star formation. The transition times of these galaxy properties are correlated with the time when the host halo transits from fast to slow accretion, indicating the two-phase assembly of halos as a potential mechanism that drives the two-phase formation of galaxies. The physical origin of dynamical hotness can be summarized into two modes of star formation: a scattered mode in which stars form at large radii within cold gas streams associated with fast assembly of halos, and a concentrated mode in which stars form at small radii through violent fragmentation from globally self-gravitated gas when halo assembly is about to slow down. Cold gaseous and stellar disks can form when the conditions of the two modes are removed by the stall of fast halo assembly and the reduction of gas by feedback processes. The two modes of star formation leave distinct imprints on the structural properties of high-redshift galaxies, providing implications to be tested by JWST and future observations.
@@
pubyear: 2026pagerange: Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos12
@@ 1 Introduction
Galaxies exhibit diverse morphologies ranging from those dominated by bulges to those dominated by disks – known as the Hubble sequence (e.g. Hubble, 1926, 1936; Sandage, 1961; Kormendy, 1979; Conselice, 2014). Observations with resolved kinematic measurements suggest that the morphologies of galaxies are closely related to the motion of stars, with bulges being populated mainly by stars with random orbits and disks with ordered rotation (e.g. Cappellari, 2016; Zhu et al., 2018). Historically, these two types of motion were inferred to reflect the evolution stages of galaxies, with ordered rotation being the early products of galaxy formation owing to the collapse of gas with initial angular momentum in halos, and random motion being generated by later ‘heating’ processes such as mergers (e.g. Hernquist, 1992; Shen et al., 2003; Lacey et al., 2016). The formation of stable disks establishes the blue, star-forming sequence. Subsequent dynamical heating and quenching of star formation, driven by merger-induced bulge growth and supermassive black hole (SMBH) feedback, cause galaxies to migrate toward the red sequence, producing the observed bimodality in the galaxy population. Such a picture of galaxy formation and morphological transformation has been widely adopted in theoretical models and achieved great success in reproducing the observed properties of the galaxy population observed in the local Universe, such as the size-mass relations (e.g. Shen et al., 2003; van Dokkum et al., 2008, 2015; Bezanson et al., 2009; Newman et al., 2012; Graham, 2013; van der Wel et al., 2014b; Somerville et al., 2018; Miller et al., 2019; Jia et al., 2024; Song et al., 2025) and the bimodality in the color-stellar mass relation (Strateva et al., 2001; Baldry et al., 2004).
@@
Along this line, Mo et al. (2024, hereafter MCW24) proposed a ‘two-phase’ scenario of galaxy formation, in which the two morphological components of galaxies, bulge and disk, naturally arise from cosmological structure formation. This scenario is motivated by the fact that the assembly of individual CDM halos follows a general pattern consisting of two phases: an early phase of fast accretion and a later phase of slow accretion – a discovery made decades ago in N-body simulations (Zhao et al. 2003; see also Wechsler et al. 2002; Zhao et al. 2009; Correa et al. 2015). During the fast phase of a halo, rapid change of gravitational potential associated with the fast accretion of the halo, complemented by energy inputs from the gas accretion itself and feedback processes, can generate large perturbations in the gas (Robertson and Goldreich, 2012; Murray et al., 2017; Fielding et al., 2017; Goldner et al., 2025; Sun et al., 2026), which can seed clouds of over-density and promote the formation of dense sub-clouds. The rich amount of gas and the strong turbulence brought in by the fast accretion imply that the fragmentation starts once the gas becomes self-gravitating and that the formation of dense sub-clouds proceeds rapidly before the gas can dissipate to form a dynamically cold disk. A galaxy formed in this phase is thus manifested as a self-gravitating gas cloud (SGC) filled with sub-clouds with random orbits, with star clusters born out of the sub-clouds inheriting their hot dynamics. Dynamically cold gaseous and stellar disk can form after the host halo transits to the slow phase and the gas content is reduced by feedback processes. Implemented semi-analytically into subhalo merger trees, this scenario can reproduce a wide range of observables, and coherently predict the co-evolution among halos, galaxies, supermassive black holes (SMBHs), and globular clusters (GCs) from the seeding era to the present day (see Mo et al., 2024; Chen et al., 2024, 2025a, 2025b, for details).
@@
One important prediction of the two-phase scenario is an order in the build-up of the Hubble sequence: the dynamically hot (bulge) component of a galaxy is actually built first, and followed by the dynamically cold (disk) component, with a transition corresponding to that in halo assembly from fast to slow. This “upside-down” order seems to be in conflict with the historical picture of galaxy formation and morphological transformation mentioned above, but becomes increasingly supported by hydrodynamic simulations (Dekel et al., 2009; Brook et al., 2012; Bird et al., 2013; Hopkins et al., 2018b; Pillepich et al., 2019; Park et al., 2019; Yu et al., 2023; Jung et al., 2025) and observations (Wuyts et al., 2012; Kassin et al., 2012; Swinbank et al., 2012; Wisnioski et al., 2015; Simons et al., 2017; Price et al., 2020). However, many questions remain unanswered. For example, it is unclear whether the fast accretion can indeed sustain a gas-rich and turbulent SGC, whether sub-clouds can actually form in the SGC quickly before settling into a rotation-supported disk, and whether feedback processes are strong enough to reduce the gas content to allow the formation of a stable disk. Given the challenges in obtaining a complete sample of high-redshift galaxies to perform dynamical decomposition, it is still difficult to answer these questions directly using observational data. On the other hand, modern zoom-in hydro simulations with cosmological initial conditions are able to resolve key processes involved in the formation of individual galaxies, thus providing opportunities to test the viability of the two-phase model.
@@
In this paper, we use a set of such simulations performed as a part of the Feedback In Realistic Environments (FIRE) project (Hopkins et al., 2018b) to analyze the co-evolution between galaxies and dark matter halos, focusing on the aforementioned questions. This paper is organized as follows. In §2, we describe the simulation and the method to quantify halos and galaxies. In §3, we show the two-phase growth pattern of galaxies, and demonstrate that it is consistent with being influenced by the two-phase assembly of dark matter halos. In §4, we analyze in detail the physical processes that drive the two-phase growth of galaxies. In §5, we provide implications for observations in a search for evidence induced by these processes. Finally, we summarize and discuss our findings in §6.
@@ 2 The Simulation
Table 1: The sample of FIRE-2 galaxies used in this paper. The properties listed in the columns are: (i) the name of the galaxy, with the reference first introducing the galaxy indicated in the parentheses (A, Garrison-Kimmel et al. 2019a; B, Garrison-Kimmel et al. 2019b; C, Garrison-Kimmel et al. 2017; D, Wetzel et al. 2016; E, Hopkins et al. 2018b; F, Samuel et al. 2020); (ii) $M_{\mathrm{vir},\mathrm{dm}}$, the mass of dark matter enclosed in the virial radius of the host halo at $z=0$; (iii) $M_{*}$, stellar mass of the galaxy; (iv) $R_{50,*}$, half-stellar-mass radius using all star particles linked to the galaxy; (v) $R_{\rm 50,bulge}$, half-stellar-mass radius using only bulge stars; (vi) $R_{90,*}$, the radius enclosing 90% of the stellar mass, using all star particles linked to the galaxy; (vii) $t_{\mathcal{B}}$, the transition time of burstiness; (viii) $t^{(\mathrm{start})}_{\mathcal{H}}$$t^{(\mathrm{end})}_{\mathcal{H}}$, the transitional interval of hotness; (ix) $^{(\mathrm{start})}_{\mathrm{thin}}tt$$^{(\mathrm{end})}_{\mathrm{thin}}$, the transitional interval of thin-disk fraction; (x) $t_{\rm bulge,1/2}$, the half-mass formation time of stellar bulge; (xi) $t^{(\mathrm{start})}_{\gamma}$$t^{(\mathrm{end})}_{\gamma}$, the transitional interval of halo accretion rate; (xii) $t_{\rm h,1/2}$, the half-mass formation time of the host halo; (xiii) $t_{\text{h, Vpeak}}$, the epoch when the maximum circular velocity of the halo reaches the peak during the assembly history. See §2 for the details of the sample and the definitions of the properties.
@@
In this paper, we use the cosmological hydrodynamic zoom-in simulations from FIRE-2, a part of the FIRE project, performed using a multimethod gravity plus (magneto)hydrodynamics code GIZMO (Hopkins, 2015), with updated numerical methods and physical models as detailed in Hopkins et al. (2018b). FIRE-2 adopted the mesh-free Lagrangian Godunov (MFM) method for hydrodynamics, which provides adaptive spatial resolution while maintaining exact conservation of mass, energy, and momentum, excellent angular momentum conservation, and accurate shock capturing. Radiative heating and cooling for gas across a temperature range of $\sim 10$$10^{10}\,{\rm K}$ are included, involving free-free, photoionization/recombination, Compton, photo-electric, metal-line, molecular, fine-structure, dust collisional, and cosmic-ray processes. Eleven elements (H, He, C, N, O, Ne, Mg, Si, S, Ca, Fe) are self-consistently tracked. Unresolved turbulent diffusion of metals in the gas is explicitly modeled (Hopkins, 2016; Su et al., 2017; Escala et al., 2018). The simulations also include photoionization and heating from a redshift-dependent, spatially uniform ultraviolet (UV) background (Faucher-Giguère et al., 2009) that reionizes the universe at $z\sim 10$. Collisionless star particles, representing ensembles of stars, are bred from molecular gas satisfying self-gravitating, self-shielding, Jeans unstable conditions and exceeding a minimum density threshold of $1000\,{\rm cm}^{-3}$ (Krumholz and Gnedin, 2011). FIRE-2 implemented distinct channels of stellar feedback, including that from OB stars, AGB mass-loss, type Ia and type II supernovae, photoelectric heating, and radiation pressure (Hopkins et al., 2018a, 2020). Such feedback regulates the star formation efficiency of star-forming gas to a level of $\sim 1\%$$10\%$ per free-fall time (Faucher-Giguère et al., 2013; Hopkins, 2017; Orr et al., 2018). We note that feedback from accreting SMBHs is not included in the simulations.
@@
Figure 2: Properties of the galaxy m12f and its host halo as functions of the cosmic time ($t$) or redshift ($z$). Each panel shows a different property. (a), stellar mass in total (gray) and separately for the three kinematic components (colored), and $M_{\rm vir}$ of the host halo (black). Three vertical lines indicate the epochs at which the bulge mass reaches 20%, 50% and 80%, respectively, of its final value at $z=0$, with that corresponding to 50% denoted as $t_{\rm bulge,1/2}$. (b), the growth rate of the three kinematic components (colored). For reference, we show the accretion rate of baryons, defined as the growth rate of $M_{\rm vir}$ times the cosmic baryon fraction (black). (c), mass fractions of the three kinematic components, using young stars only. Grey vertical band indicates the transitional interval, $_{\rm thin}^{\rm(start)}tt$$_{\rm thin}^{\rm(end)}$, of the thin-disk fraction. (d), dynamical hotness of the galaxy ($\mathcal{H}$), calculated using young stars only. Grey vertical band indicates the transitional interval, $t_{\mathcal{H}}^{\rm(start)}$$t_{\mathcal{H}}^{\rm(end)}$ of $\mathcal{H}$. (e), burstiness of star formation ($\mathcal{B}$). Vertical line indicates $t_{\mathcal{B}}$, the epoch when the star formation transits from bursty to steady (defined by $\mathcal{B}=0.2$, indicated by the dashed horizontal line). (f), cold-gas spin $\lambda_{\text{cold-gas}}$, and fractions of cold ($f_{\text{cold-gas}}$), warm and hot gas. (g), (h), virial velocity ($V_{\rm vir}$) and specific growth rate of the host halo ($\gamma$; see Eq. 15 for definition). Grey vertical bands indicate the transitional interval, $t_{\gamma}^{\rm(start)}$$t_{\gamma}^{\rm(end)}$, of halo assembly, defined by $\gamma=3/8$0 (horizontal lines in h) In (c), (d), (g) and (h), each solid curve is obtained from the simulation, while the associated dashed curve is obtained by a parametric fitting (see §2.1 and §2.2 for details). This figure shows that the galaxy evolves from an early phase featured by hot dynamics, bulge-dominated structure and bursty star formation, to a later phase featured by cold dynamics, disk-dominated structure and steady star formation, and that the assembly of the host halo decelerates with time. See §3 for details.
@@ 2.1 Properties of the simulated galaxies
$\epsilon$ Orbital circularity of a particle (§2.1.1)
@@ 2.1.1 Stellar kinematics and decomposition
Kinematic decomposition in this paper is performed using a circularity-based method similar to that of Yu et al. (2021, 2023). For each galaxy at a given snapshot, we categorize each of its member star particles into one of the three components: bulge, thick disk and thin disk, according to the circularity (Abadi et al., 2003) of the star particle:
@@
where $j_{z}$ is the angular momentum of that particle projected to the direction of $\bm{J}_{\rm tot}$, the net angular momentum of all stellar particles within $R_{\rm 90,*}$ at that snapshot (Gurvich et al., 2023b), $j_{\rm c}(E)$ is the angular momentum of a circular orbit with the same energy ($E$) as that of the star particle. Particles with $1\leqslant\epsilon<0.2$, $0.2\leqslant\epsilon<0.8$ and $0.8\leqslant\epsilon\leqslant 1$ are categorized as bulge, thick disk and thin disk, respectively. Previous studies adopted more rigorous kinematic classification methods (e.g. Du et al., 2019, 2020; Zana et al., 2022; Liang et al., 2025), such as classifications in the plane of angular momentum versus orbital energy. However, our work focuses primarily on morphological transitions, particularly in the evolution of the significance of the disk component, which can be characterized robustly using the angular momentum alone. We therefore employ a simplified kinematic decomposition based solely on the angular momentum to define the three components. We define the direction of $\bm{J}_{\rm tot}$ as the normal direction (the $z$ axis) of the disk ($x$-$y$) plane. The projections of a simulated galaxy perpendicular to and along the $z$-axis are defined as the edge-on and face-on views of the galaxy, respectively.
@@
Fig. 1 shows an example galaxy, m12f, at $z=0$, with the three components shown separately and in total, viewed edge-on and face-on. Here, we include all particles within $R_{\rm 90,*}$. Panel (e) shows the distribution of $\epsilon$ for all stellar particles within $R_{\rm 90,*}$, weighted by their masses. Panel (f) shows the face-on surface density profile of the three components and of the total, using all particles within $R_{\rm 90,*}$. The thin disk is extended, dominates the surface density in the outer region, and follows an exponential profile at $r\gtrsim 3\,{\rm{kpc}}\approx R_{\rm 50,*}$. In the inner region ($r\lesssim 1\,{\rm{kpc}}$), however, the thin disk is truncated. This truncation is similar to that found in other simulations (Du et al. 2020, see their fig. 1; Buck et al. 2019, see their fig. 2) and in observations with spatially-resolved spectroscopy (Zhu et al. 2018, see their fig. 8 and §4), and suggests a potential caveat of using an un-truncated profile to fit the entire cold-disk component. On the other hand, the thick disc is more vertically extended, reflecting its hotter dynamics with a higher degree of random-motion support for the vertical structure. Both the images (c1, c2) and the surface density profile (yellow in f) suggest the existence of an inner, concentrated component at around $1\,{\rm{kpc}}$, in alignment with the concept of slow-rotating “pseudo-bulge” found in observations (e.g. Kormendy and Kennicutt, 2004; Fisher and Drory, 2016; Hu et al., 2024) and in hydrodynamic simulations (e.g. Okamoto, 2013; Gargiulo et al., 2019; Du et al., 2020; Zana et al., 2022). The bulge dominates the inner $0.5\,{\rm{kpc}}$, but it also has a faint extension to the outer region ($\gtrsim 10\,{\rm{kpc}}$), consistent with being a combination of the “classical bulge” and the stellar halo found in observations (e.g. Kormendy and Kennicutt, 2004; Fisher and Drory, 2016; Huang et al., 2018a; Hu et al., 2024) and in hydrodynamic simulations (e.g. Du et al., 2020; Zana et al., 2022).
The three components have similar mass within $R_{\rm 90,*}$, as can be seen from the mass-weighted distribution of $\epsilon$ (panel e), consistent with the results of an earlier analysis of the FIRE-2 by Yu et al. (2021). All these show that FIRE-2 is able to produce realistic kinematics of MW-size galaxies, with components similar to those found in observations.
@@
To quantify the kinematics of star-forming gas within a target galaxy, we utilize in-situ young stars, defined as bound stellar particles with ages $\leq 100\,{\rm Myr}$ and birthplaces within $\leq R_{90,*}$ from the center of the galaxy. This selection is made to avoid complexity introduced by potential post-formation heating of old stars and the contamination by ex-situ stars. Hereafter, we use these in-situ young stars to define transition times in the morphological type of galaxy. As shown in Fig. 1(e), the distribution of $\epsilon$ of in-situ young stars at $z=0$ is sharply peaked around $\epsilon=1$. Most ($\approx 93\%$) of the in-situ young stars at this epoch belong to the thin disk, and only a small fraction ($\approx 1.6\%$) is categorized into the bulge component. Fig. 1(f) reinforces this conclusion by comparing the face-on surface densities of in-situ young stars (dashed lines) with those of all stars (solid lines), which shows that in-situ young stars are mostly contained in the extended thin disk.
@@
The transition of star formation from the dynamically hot to cold components shown in Fig. 2(c) motivates us to define a transition time to separate the early, hot phase from the late, cold phase, for each individual galaxy. To do this, we fit the fraction of young stars in the thin-disk component ($f_{\rm thin}$) as a function of cosmic time ($t$), as shown by the blue solid curve in Fig. 2(c) for example, by the following piecewise function: